Determine the energy of a photon with a wavelength of 490nm
Answer #1

Formula EF nc Wavelength, x=490, nm Planet constant, h = 6-426X1634 Is speed of light, C = 3x1678 m/s -34 E = 6-626x 10 x 3 x